# From Conversation to Automation: Leveraging Large Language Models to Analyze Strategies in Problem Solving Therapy

### v1

@@ 4.1 Performance of LLMs in Annotation
Table 1 shows the weighted F1 scores of four LLMs, including two proprietary models (GPT-4 and GPT-4o) and two open-source models (Llama-3.1 and Yi-1.5). We evaluate the models’ performance using our evaluation set (500 annotated utterances) for both ‘No context’ and ‘With context’ conditions. GPT-4o achieved the highest F1 score (0.76) with ‘No context’, outperforming all other models. GPT-4 followed closely with F1 score of 0.68. All models showed lower F1 scores when provided with additional context (i.e., the two previous utterances in the conversation). For GPT-4, the performance decreased from 0.68 to 0.64 when context was included. GPT-4o experienced a more significant drop, from 0.76 to 0.61. Similarly, for the smaller open-source models, Llama-3.1 and Yi-1.5, performance dropped to F1 scores of 0.50 (See Appendix E.1 for more details).
Table 2 summarizes the entropy results across five runs for each LLM. Since entropy values were the same for both ‘No context’ and ‘With context,’ we report only one set of results. The entropy values across all models were low enough to ensure consistency.Both GPT-4 and GPT-4o had higher mean entropy than Llama-3.1 and Yi-1.5, with GPT-4o slightly more consistent (0.035 vs. 0.042). Although GPT models showed higher variability (standard deviations of 0.150 and 0.130), their higher F1 scores indicate this uncertainty had no impact on the overall performance of the models. In contrast, Llama-3.1 and Yi-1.5 had much lower entropy, with Yi-1.5 showing complete consistency (0.00), but this may reflect overconfidence and poor generalization, as suggested by the lower F1 score.
Overall, the classification result shows that GPT-4o performed best without any context. Precision, recall, and F1 scores for each PST strategy are shown in Table 3. The model performed well in classifying ‘PS core’ strategies, with balanced scores across strategies like ‘Problem-Solving Positive Mindset’ and ‘Defining Problems and Goals’. For communication, it excelled in identifying ‘Therapeutic Engagement’ and ‘Problem-Solving Test Review’, while ‘Social Courtesies’ and ‘Session Management’ showed slightly lower precision. As a result of these experiments, we selected GPT-4o to annotate the entire dataset of therapist utterances. Given its low mean entropy (0.035), we only ran the model once for full annotation.
@@
GPT4 0.68 0.64
GPT4o 0.76 0.61
Llama-3.1 0.61 0.50
Yi-1.5 0.56 0.50
Table 1: F1 Score (Weighted Average) Comparison Between LLMs
@@
Table 2: Comparison of Mean and Standard Deviation of Entropy across LLMs
@@
Communication Social Courtesies 0.59 0.94 0.73 25
@@
Problem-Solving Test Review 1 0.86 0.92 24
@@
Table 3: Precision, Recall, and F1-Score for Communication and PS Strategies in GPT4o (without context)
4.2 Model Fine-Tuning Results
After annotating the entire dataset of 14,417 therapist utterances with GPT-4o, we randomly sampled 5,000 utterances and fine-tuned three transformed-based models. We used the 500 utterances annotated by human experts as the evaluation set. Table 4 shows the average weighted F1 scores for DeBERTa, MentalBERT, and FLAN-T5 on ‘PS core’ and ‘Communication’ strategies. MentalBERT achieved the highest F1 score (0.78) for ‘PS core’ strategies, while DeBERTa performed best on ‘Communication’ (0.73). Both models tied for overall performance with an F1 score of 0.68, while FLAN-T5 slightly trailed at 0.67. These results highlight DeBERTa’s strength in ‘Communication’ annotation and MentalBERT’s advantage in ‘PS core’ (See Appendix E.2).
Model PS Commu. Combined
@@ 4.2 Model Fine-Tuning Results
Table 4: Average weighted F1 scores for PS Core (PS), Communication (Commu.), and overall across models.
@@
Communication Session Management (next, time), (time, meet), (last, time), (defining, problem), (alternative, solutions), (next, week), (let, see), (going, talk), (time, talk), (brainstorming, ideas) focusfuture 15.80%(1,792)
@@
Problem-Solving Test Review (problem, solving), (positive, rational), (impulsive, avoidant), (positive, negative), (high, positive), (rational, impulsive), (negative, orientation), (negative, rational), (approach, problem), (one, questionnaires), (positive, orientation), (five, dimensions) insight 1.72%(196)
@@
4.3 Characteristics of the PST Data
@@
Language Patterns in Therapeutic Strategies. Table 5 shows the distribution of therapeutic strategies and top bigrams in our labeled dataset by GPT4o. To gain deeper insights into the linguistic features of each strategy, we further analyzed the labeled utterances using the Linguistic Inquiry and Word Count (LIWC) Boyd et al. (2022). Among the ‘PS core’ strategies, ‘Defining Problems and Goals’ is the most prevalent (33.48%). This highlights the therapists’ emphasis on helping clients clarify their objectives, with words like “defining problem” and “solving important”. This strategy is further supported by LIWC features such as ‘reward,’ suggesting clients are motivated by the potential benefits or outcomes tied to their goals. ‘Generating Alternative Solutions’ is another significant strategy (20.03%), with words such as “brainstorming ideas” and “strategies tactics” showing the focus on exploring different options to address issues. The LIWC indicators ‘insight’ and ‘curiosity’ align with this focus on creativity and open-mindedness. Similarly, ‘Problem-Solving Positive Mindset’ bigrams, such as “positive attitude” and “healthy thinking”, show therapists’ encouragement to adopt a more constructive outlook on problems, fostering optimism. The prevalence of ‘emotion’ and ‘feeling’ reflects the emotional engagement involved in developing such positive mindset. ‘Outcome Prediction and Planning’ phrases like “pros cons” and “solve problem” indicate a focus on helping clients think through potential consequences. This is echoed by the prevalence of ‘risk’ and ‘focus future’ as dominant features, reinforcing the importance of careful consideration and future-oriented thinking.
@@
On the ‘Communication’ side, the bigrams highlight how therapists build rapport and guide session flow. In ‘Therapeutic Engagement’, phrases like “really good” use affirming language to foster a supportive, collaborative atmosphere, helping clients feel understood. ‘Social Courtesies,’ with phrases like ‘thank much’ and ‘look forward,’ add warmth to interactions, maintaining a respectful and friendly tone. The top LIWC features, ‘emotion’ and ‘prosocial,’ reflect the positive social behaviors therapists use to foster mutual respect and connection. Meanwhile, in ‘Problem-Solving Test Review’, more formal words like “positive rational” and “impulsive avoidant” suggest an analytical approach, where therapists review the client’s PST test results in a structured, evaluative manner (See Appendix F for more information).
@@
Progression of Strategies Across Therapy Visits. Figures 3 and 3 show a clear progression of strategies across three different visits of caregivers. In Figure 3, early sessions focus heavily on enhancing ‘Problem-Solving Mindset’ (Step One) and ‘Defining Problems and Goals’ (Step Two), indicating that the initial phase of therapy is centered on improving caregivers’ mindsets and helping them clearly define the challenges they face. As therapy progresses into the second and third visits, there is a shift towards more advanced strategies such as ‘Generating Alternative Solutions’ (Step Three) and ‘Outcome Prediction and Planning’ (Step Four), showing that the focus moves towards exploring possible solutions and planning for their implementation. By the third visit, ‘Trying Out Solution Plan’ (Step Five) becomes more prominent, suggesting that the final phase of therapy is geared towards applying and testing these solutions in real-life situations. Figure 3 shows the distribution of ‘Communication’ strategies across visits. Strategies such as ‘Therapeutic Engagement’, ‘Session Management’, and ‘Social Courtesies’, were applied consistently throughout the sessions. This indicates that these strategies are essential components of every visit, regardless of the stage of therapy. However, ‘PS Test Review’ was used more frequently during the first visit, reflecting the therapist’s focus on reviewing the client’s Problem-Solving Test results, which were completed before the therapy began. This early emphasis helps set the foundation for the sessions that follow. These intuitive results support the effectiveness of our classifier.
@@ 5 Discussion
Annotating PST-Related Texts. Our results show that GPT-4o significantly outperformed both GPT-4 and various open-weight models in annotating therapeutic strategies, achieving F1 score of 0.76 without the need for supplemental context. Despite Yi exhibiting the lowest entropy among the models tested, GPT-4o displayed higher consistency and reliability. This highlights critical trade-offs to consider between performance consistency and overall model reliability when deploying LLMs in sensitive fields such as mental health care. While GPT-4o demonstrates strong effectiveness, its success highlights the need for advancements in open-source models. Enhancing these models’ accuracy and reliability is essential to providing competitive and accessible alternatives in the healthcare domain. Such advancements could democratize access to high-quality mental health resources, thereby broadening the impact and reach of therapeutic interventions. Our models overall show better results compared to previous work Chiu et al. (2024), which reported that their best-performing classifier, GPT-4 with definitions and examples, achieved an F1 score of 57.7 in identifying therapist behavior across 13 different therapist codes.
Figure 2: Distribution of PS core Strategies in Different Visits
Figure 3: Distribution of Communication Strategies in Different Visits
Comparing the performance of the models in labeling ‘PS core’ vs. ‘Communication’ strategies, GPT-4o demonstrated notable efficacy in labeling structured, goal-oriented ‘PS core’ strategies such as ‘Defining Problems and Goals’ and ‘Generating Alternative Solutions,’ achieving high precision and recall. This performance highlights LLMs’ strength in recognizing structured, task-oriented language that is clear and objective.
On the other hand, GPT-4o was less robust when dealing with ‘Communication’ strategies like ‘Session Management’ and ‘Social Courtesies.’ These strategies are more dynamic and rely heavily on emotional and social cues. This disparity in performance highlights a key challenge for LLMs in therapeutic settings: while they excel at handling clear, problem-solving tasks, they struggle with the subtleties of interpersonal communication essential in therapy, lacking the ability to accurately interpret the nuances of human emotions and social interactions. These characteristics can often be implicit and vary significantly across different individuals and situations. Enhancing LLMs’ sensitivity to these subtle dynamics is important for improving their effectiveness within therapeutic environments. Research has indicated that LLMs can sometimes exhibit an over-realistic sense of empathy Sorin et al. (2023), stemming from a fundamental misunderstanding of underlying emotional contexts. To address this issue, future developments should concentrate on refining LLMs by incorporating strategies such as expert-in-the-loop systems Afzal et al. (2024); Sun et al. (2023). These systems should allow for real-time feedback and guidance from human experts, which could enhance the models’ ability to recognize and interpret complex emotional and social cues.
Model Fine-Tuning and Generalizability. The results of our fine-tuned transformer-based models showed that MentalBERT performed best in annotating problem-solving core strategies with an F1 score of 0.78, while DeBERTa excelled in communication strategies with a 0.73 F1 score. This indicates that different models can specialize in distinct aspects of therapeutic dialogue. Interestingly, our transformer-based model achieved an overall F1 score of 0.68, comparable to our second-best LLM.This result is particularly noteworthy as it suggests that both MentalBERT and DeBERTa, with their distinctive capabilities, can serve as reliable and accessible annotators for a range of therapeutic strategies, presenting a promising avenue for scalable and efficient analysis of therapy content. These models can also be more appropriate in the context of therapy where private and sensitive information is being analyzed Shah et al. (2022). This capability is important for maintaining confidentiality and trust in digital therapeutic environments, which is fundamental for effective therapy.
Real-World Applications of LLM-Annotated Data. The strong performance of our models, especially GPT-4o, in annotating therapeutic strategies, even without conversational context, demonstrates the potential for using LLMs to automate the analysis of PST transcripts at scale. By integrating LLMs into real-time therapy support or documentation systems, therapists could receive automated insights that help them refine their approaches, track client progress, and identify appropriate strategies more efficiently (As shown in Figures 3 and 3). These systems could also tailor interventions to individual client needs and offer real-time suggestions to therapists, further optimizing therapeutic outcomes. However, this raises important questions about the risks of over-reliance on AI-generated feedback. Our findings indicate that while LLMs are adept at processing and analyzing data, they sometimes fail to accurately identify specific therapeutic strategies and lack the capability to convey genuine empathy and emotion. While LLMs can provide valuable assistance, ensuring that human judgment remains central to the therapeutic process is crucial, as LLMs may struggle with the nuanced emotional and interpersonal dynamics that are critical in therapy. Balancing AI support with therapist oversight will be key to maximizing the benefits of these technologies while minimizing potential risks.

### v2

@@ 4.1 Performance of LLMs in Annotation
Table 1 shows the weighted F1 scores of four LLMs, including two proprietary models (GPT-4 and GPT-4o) and two open-source models (Llama-3.1 and Yi-1.5). We evaluate the models’ performance using our evaluation set (500 annotated utterances) for both ‘No context’ and ‘With context’ conditions. GPT-4o achieved the highest F1 score (0.76) with ‘No context’, outperforming all other models. GPT-4 followed closely with F1 score of 0.68 (See Appendix E.1 for more details). All models showed lower F1 scores when provided with additional context (i.e., the two previous utterances in the conversation), and our error analysis revealed that this was because, despite explicitly instructing the LLMs in our prompt to focus only on the therapist’s most recent utterance and use the two previous utterances solely as context, the LLMs occasionally returned labels for the context rather than the target utterance.
Table 2 shows entropy results across five runs per LLM. Since values were identical for ‘No context’ and ‘With context’, we report a single set, with all models showing sufficiently low entropy for consistency. GPT-4 and GPT-4o had higher mean entropy than Llama-3.1 and Yi-1.5, with GPT-4o slightly more stable (0.035 vs. 0.042). Despite higher variability, GPT models’ superior F1 scores suggest this had no impact on performance. In contrast, Llama-3.1 and Yi-1.5 exhibited lower entropy, with Yi-1.5 at 0.00, potentially indicating overconfidence and poor generalization, reflected in its lower F1 score.
Overall, the results show that GPT-4o performed best without any context. Precision, recall, and F1 scores for each PST strategy are shown in Table 3. The model performed well in classifying PS Core strategies, with balanced scores across strategies like ‘Problem-Solving Positive Mindset’ and ‘Defining Problems and Goals’. For Facilitators, it excelled in identifying ‘Therapeutic Engagement’ and ‘Problem-Solving Test Review’, while ‘Social Courtesies’ and ‘Session Management’ showed slightly lower precision. We selected GPT-4o to annotate the entire dataset of therapist utterances. Given its low mean entropy (0.035), we only ran the model once for full annotation.
@@
FAC PS Overall FAC PS Overall
GPT-4 0.65 0.76 0.68 0.65 0.70 0.64
GPT-4o 0.78 0.81 0.76 0.59 0.76 0.66
Llama-3.1 0.66 0.70 0.61 0.55 0.65 0.50
Yi-1.5 0.64 0.59 0.56 0.56 0.54 0.49
Table 1: Average weighted F1 scores for PS Core (PS), Facilitators (FAC), and overall across different LLMs. The overall score is the weighted average of all strategies’ F1 score plus None category.
@@
Table 2: Comparison of Mean and Standard Deviation of Entropy across different LLMs.
@@
Facilitators Social Courtesies 0.59 0.94 0.73 25
@@
Test Review 1 0.86 0.92 24
@@
Table 3: Precision, Recall, and F1-Score for Facilitative and PS Strategies in GPT-4o (without context)
4.2 Transformer-Based Models
Table 4 shows the average weighted F1 scores for DeBERTa, MentalBERT, FLAN-T5, and ModernBERT on PS Core and Facilitative strategies. ModernBERT achieved the highest F1 score (0.8) for PS Core strategies, while DeBERTa performed best on Facilitators (0.73). These results highlight DeBERTa’s strength in Facilitators annotation and ModernBERT’s advantage in PS Core (See Appendix E.2).
Model PS FAC Overall
@@ 4.2 Transformer-Based Models
Table 4: Average weighted F1 scores for PS Core (PS), Facilitators (FAC), and Overall across models.
@@
Facilitators Session Management (next, time), (time, meet), (last, time), (defining, problem), (alternative, solutions), (next, week), (let, see), (going, talk), (time, talk), (brainstorming, ideas) focusfuture 15.80%(1,792)
@@
Test Review (problem, solving), (positive, rational), (impulsive, avoidant), (positive, negative), (high, positive), (rational, impulsive), (negative, orientation), (negative, rational), (approach, problem), (one, questionnaires), (positive, orientation), (five, dimensions) insight 1.72%(196)
@@
Figure 2: Distribution of Autonomy in PS Core
Figure 3: Distribution of Autonomy in Facilitators
4.3 Therapeutic Dynamics
For autonomy, the relationship between directive and non-directive language and question types shows that 94.49% of directive utterances are closed-ended. while 59.25% of non-directive utterances are open-ended, a notable 40.75% are unexpectedly closed-ended (See the co-occurrence matrix in Appendix G). Figure 3 shows the distribution of directive and non-directive autonomy in PS Core strategies. Non-directive language dominates the early stages (‘PS Positive Mindset’, ‘Defining Problems and Goals’), while the later stages (‘Outcome Prediction and Planning’, ‘Trying Out Solution Plan’) shift toward directive language. Similarly, for Facilitators (Figure 3) more structured strategies are directive while the other two show more non-directive language.
Reviewing the results of self-disclosure showed that, only 257 instances were labeled as non-immediate self-disclosure, suggesting that such disclosures are carefully considered rather than spontaneous. Furthermore, Figures 5 and 5 show metaphor usage across PS Core and Facilitative Strategies. Metaphors are most prevalent in Therapeutic Engagement (72.18%). In PS Core Strategies, they are commonly used in ‘Defining Problems and Goals’ (44.90%) and ‘Generating Alternative Solutions’ (23.13%).
4.4 Linguistic Patterns
@@
Table 5 shows the distribution of all therapeutic strategies in our data with the top bigrams, and LIWC categories representing each. Among PS Core strategies, ‘Defining Problems and Goals’ is most prevalent (33.48%), emphasizing therapists’ role in helping clients clarify objectives, with bigrams like “defining problem” and “solving important”. LIWC’s ‘reward’ feature suggests clients are being motivated by potential benefits. ‘Generating Alternative Solutions’ is another significant strategy (20.03%), with words such as “brainstorming ideas” and “strategies tactics” showing the focus on exploring different options to address issues. The LIWC indicators ‘insight’ and ‘curiosity’ align with this focus on creativity and open-mindedness. Similarly, ‘Problem-Solving Positive Mindset’ includes “positive attitude” and “healthy thinking”, highlighting therapists’ encouragement toward optimism and resilience, supported by LIWC features ‘emotion’ and ‘feeling’. Outcome Prediction and Planning uses “pros cons” and “solve problem”, emphasizing consequence evaluation. The prevalence of LIWC’s ‘risk’ and ‘focus future’ reinforces a forward-thinking approach in decision-making.
@@
On the ‘Facilitators’ side, the bigrams highlight how therapists build rapport and guide session flow. In ‘Therapeutic Engagement’, phrases like “really good” use affirming language to foster a supportive, collaborative atmosphere, helping clients feel understood. ‘Social Courtesies,’ with phrases like ‘thank much’ and ‘look forward,’ add warmth to interactions, maintaining a respectful and friendly tone. The top LIWC features, ‘emotion’ and ‘prosocial,’ reflect the positive social behaviors therapists use to foster mutual respect and connection. Meanwhile, in ‘Test Review’, more formal words like “positive rational” and “impulsive avoidant” suggest an analytical approach, where therapists review the client’s PST test results in a structured, evaluative manner (See Appendix F for more information).
4.5 Strategic Progression in Therapy Sessions.
@@
We analyzed how strategies progress across three different visits of caregivers in our dataset. In Figure 7, early sessions focus on enhancing ‘Problem-Solving Mindset’ (Step One) and ‘Defining Problems and Goals’ (Step Two), indicating that the initial phase of therapy is centered on improving caregivers’ mindsets and helping them define the challenges they face. As therapy progresses into the second and third visits, there is a shift towards more advanced strategies such as ‘Generating Alternative Solutions’ (Step Three) and ‘Outcome Prediction and Planning’ (Step Four), showing that the focus moves towards exploring possible solutions and planning for their implementation. In the final phase (Visit Three), ‘Trying Out Solution Plan’ (Step Five) becomes prominent, indicating a transition to applying and testing solutions in real life. Figure 7 shows the consistent use of Facilitative strategies like ‘Therapeutic Engagement’, ‘Session Management’, and ‘Social Courtesies’, reinforcing their essential role throughout therapy. However, ‘Test Review’ is more frequent in Visit One, reflecting the therapist’s focus on reviewing the client’s Test results taken prior to starting the session before progressing to intervention. This early emphasis lays the groundwork for later sessions. These patterns further validate the effectiveness of our classifier.
Figure 4: Proportion of Metaphors in PS Core
Figure 5: Proportion of Metaphors in Facilitators
@@ 5 Discussion
Model Performance and the Need for Open-Source Improvements. Our results show that GPT-4o outperformed GPT-4 and open-weight models in annotating therapeutic strategies, achieving an F1 score of 0.76 without supplemental context. While Yi had the lowest entropy, GPT-4o showed greater consistency and reliability, highlighting the trade-off between performance stability and model reliability in mental health applications. Despite GPT-4o’s success, improving open-source models is crucial for enhancing accuracy, accessibility, and competition in healthcare AI. From an ethical perspective, open-source models promote transparency and accountability, ensuring that healthcare AI advancements benefit a broader community and expand access to high-quality mental health resources—unlike proprietary models, which can limit access and obscure critical decision-making processes Welsh et al. (2024).
Figure 6: Distribution of PS Core in three visits
Figure 7: Distribution of Facilitators in three visits
Improving LLMs for Emotional and Social Understanding. GPT-4o excelled in labeling structured, goal-oriented PS core strategies, achieving high precision and recall. This highlights LLMs’ strength in processing clear, task-oriented language. However, GPT-4o struggled with Facilitative strategies such as ‘Session Management’ and ‘Social Courtesies’, which rely on emotional and social cues. This underscores a key limitation: while LLMs handle problem-solving tasks well, they struggle with the nuances of interpersonal communication, lacking the ability to interpret human emotions accurately. These challenges arise from implicit, context-dependent interactions, which LLMs often misinterpret, sometimes displaying an over-realistic sense of empathy Sorin et al. (2023). To improve effectiveness in therapeutic settings, future work should refine LLMs using expert-in-the-loop systems, allowing for real-time human feedback to enhance emotional and social cue recognition Afzal et al. (2024); Sun et al. (2023).
Efficient and Privacy-Conscious Therapy Annotation. Our fine-tuned transformer-based models show the effectiveness of the distillation (student-teacher) method, achieving strong performance. ModernBERT outperformed our second-best LLM, GPT-4, with an overall F1 score of 0.71. It achieved 0.8 F1 in PS Core strategies, while DeBERTa specialized in Facilitative strategies, reaching an F1 score of 0.73. This suggests that smaller models can effectively annotate distinct therapeutic strategies, offering a scalable and efficient alternative for therapy content analysis. Additionally, these models provide a privacy-conscious solution for processing sensitive therapy data, ensuring confidentiality and trust in digital therapeutic environments Shah et al. (2022).
How PST is Delivered. Beyond classifying strategies, understanding how therapists apply therapeutic tools like autonomy, self-disclosure, and metaphor offers deeper insight into PST’s communication dynamics. In early PS Core stages, therapists favor non-directive language to foster client exploration, shifting to directive language later for clarity and structure, aligning with a client-centered problem-solving approach Nezu et al. (2012). Examining question types and autonomy, we found that some closed-ended questions still encouraged elaboration, which our domain expert termed “imposters.” For example, “Do you have any questions about what we’ve gone over or for this process?”—though closed-ended, it invites further discussion, blurring the line between directive and non-directive communication. This suggests that therapists sometimes frame questions strategically to encourage deeper engagement, even when using traditionally restrictive formats.
Therapists also use non-immediate self-disclosure strategically and selectively, as reflected in the low frequency of such utterances in our dataset. In Facilitators, self-disclosures often involve personal anecdotes to build rapport, whereas, in PS Core, they serve a more functional role, modeling problem-solving strategies for clients. This finding is supported by previous work Knox and Hill (2003) which identified similar patterns of strategic disclosure timing and emphasized how self-disclosure can serve as different functions at different stages of therapy. Metaphors help illustrate ideas and enhance emotional connection in Therapeutic Engagement Lyddon et al. (2001); Tay (2014), while in PS Core strategies, they help conceptualization and creativity. However, they decline in action-oriented phases, where direct communication becomes essential. Therapists adapt their language to session goals—fostering exploration early on and shifting to structured, goal-driven communication as therapy progresses.
Broader Implications. Our model has meaningful real-world applications for PST practice and documentation. It can assist annotating and analyzing therapy sessions, helping document which PST strategies were employed and how effectively they were implemented, and serving valuable educational purposes for therapists-in-training. As LLM-assisted therapy models become more prevalent, our work provides important guidance for human-AI collaborative approaches. By identifying which PST strategies are most crucial and how they should be properly sequenced for building client rapport, we can better inform the development of LLM-assisted therapy systems where human therapists remain in the loop. The model can also serve as an evaluation framework for assessing whether AI-assisted therapy is appropriately implementing PST strategies and maintaining therapeutic alliances through their interactions.